### v1

@@ V-A Uncertainty-aware Softmax
Conventional Softmax-based objectives (e.g., Eq. 7 and Eq. 9) operate solely on the deterministic embedding $\boldsymbol{\phi}^{s}$, which represents only a point estimate of the speaker identity. Such formulations overlook the covariance information $\mathbf{\Sigma}^{s}$, thereby discarding valuable information about embedding reliability.
@@
To address this limitation, we reformulate the normalization equation 8 by incorporating the uncertainty term, the uncertainty-aware normalized embedding is given by:
$\tilde{\boldsymbol{\phi}}^{s}$ $=\frac{\sqrt{{\boldsymbol{\phi}^{s}}^{\top}(\mathbf{\Lambda}+\mathbf{\Sigma}^{s})^{-1}\boldsymbol{\phi}^{s}}}{\|\boldsymbol{\phi}^{s}\|},$ (15)
which normalizes the embedding under the Mahalanobis metric induced by the uncertainty covariance $\mathbf{\Lambda}+\mathbf{\Sigma}^{s}$ where $\mathbf{\Lambda}$ is a bias term that enhances numerical stability to avoid small denominators. This formulation ensures that $\tilde{\boldsymbol{\phi}}^{s}$ is a unit vector in the uncertainty-aware space, reflecting both its magnitude and its uncertainty structure. However, in practice, explicitly computing the inverse $(\mathbf{\Lambda}+\mathbf{\Sigma}^{s})^{-1}$ often leads to unstable gradients or NaN values during training. To mitigate this issue, we adopt a numerically stable alternative:
$\tilde{\boldsymbol{\phi}}^{s}$ $=\frac{\|\boldsymbol{\phi}^{s}\|}{\sqrt{{\boldsymbol{\phi}^{s}}^{\top}(\mathbf{\Lambda}+\mathbf{\Sigma}^{s})\boldsymbol{\phi}^{s}}}.$ (16)
These two expressions are closely related. We visualize their values in Fig. 2, where both the bias term $\Lambda$ and the embeddings $\boldsymbol{\phi}$ are generated randomly. The figure demonstrates that the trends of the two equations are similar, with the primary difference being their absolute values. Moreover, as the mean variance increases, the discrepancy between the two expressions also tends to grow.
Figure 2: The value difference between Equation 15 and Equation 16.
Furthermore, this second, stable expression can be decomposed into two distinct components:
@@
$\tilde{\boldsymbol{\phi}}^{s}$ $=\frac{\boldsymbol{\phi}^{s}}{\|\boldsymbol{\phi}^{s}\|}\cdot\frac{\|\boldsymbol{\phi}^{s}\|}{\sqrt{{\boldsymbol{\phi}^{s}}^{\top}(\mathbf{\Lambda}+\mathbf{\Sigma}^{s})\boldsymbol{\phi}^{s}}},$ (17)
$=\hat{\boldsymbol{\phi}}^{s}\cdot s_{u},$ (18)
@@
where $\hat{\boldsymbol{\phi}}^{s}$ is a unit-norm embedding and $s_{u}>0$ is referred to as the uncertainty-aware scale. Crucially, this scale increases as the uncertainty decreases, effectively modulating the embedding’s magnitude based on the uncertainty of its estimate. We then incorporate the uncertainty-aware embedding $\tilde{\boldsymbol{\phi}}^{s}$ into the AAM-Softmax formulation to obtain the uncertainty-aware AAM-Softmax ($\text{AAM}_{u}$):
$\mathcal{L}_{\text{AAM}_{u}}=$ $-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{s\cdot s_{u}\cos(\theta_{y_{i}}+m)}}{e^{s\cdot s_{u}\cos(\theta_{y_{i}}+m)}+\sum_{j=1,j\neq y_{i}}^{C}e^{s\cdot s_{u}\cos\theta_{j}}}.$ (19)
@@ V-B Gradient-Driven Adaptation of Scale and Covariance
From a training perspective, the uncertainty-aware scale $s_{u}$ allows the model to adjust its confidence dynamically based on the reliability of each embedding. For an unreliable sample, typically corresponding to a large loss, the optimizer naturally encourages the model to reduce scale during back-propagation [56, 52]. A smaller scale softens the softmax distribution, reducing the penalty from incorrect or ambiguous logits and preventing the model from making overly confident but potentially wrong predictions. In our design, decreasing $s_{u}$ directly corresponds to increasing the predicted variance (uncertainty), which aligns with intuitive expectations: harder samples should be associated with higher uncertainty.
Conversely, for reliable samples with small loss, the gradient pushes $s_{u}$ upward. This sharpens the softmax distribution, enhancing the separation between target and non-target classes, and leads to smaller predicted uncertainty. In this way, the proposed scale-variance mechanism provides a training-driven confidence adjustment: for hard samples, which correspond to large losses, the back-propagation gradient drives the model to decrease the uncertainty-aware scale $s_{u}$. As a result, the predicted variance $\mathbf{\Sigma}^{s}$ increases, reflecting the higher uncertainty of these embeddings. Conversely, for easy samples with small loss, the gradient encourages $s_{u}$ to increase, leading to smaller $\mathbf{\Sigma}^{s}$, as shown in Fig. 3.
Hard Sample(Large Loss)Backpropagation GradientDecrease $s_{u}$Increase $\mathbf{\Sigma}^{s}$Easy Sample(Small Loss)Backpropagation GradientIncrease $s_{u}$Decrease $\mathbf{\Sigma}^{s}$
Figure 3: Illustration of how the uncertainty adjusts based on sample difficulty.
@@ VI Uncertainty Estimation Module
The previous two sections focus on modifying the training strategies without altering the model architecture. In this section, we aim to redesign the uncertainty estimation module. Our previous work xi-vector [24], the uncertainty estimation relies on two simple learnable layers with basic activation functions to predict frame-level uncertainty, as illustrated in Fig. 1. However, this design is relatively shallow and lacks the ability to capture temporal dependencies across speech frames. Our previous work [36] addressed this limitation by incorporating a Transformer encoder to explicitly model temporal relations, thereby enhancing the model’s capacity to capture complex non-linear patterns. Although Transformer-based architectures have been explored for speaker recognition, they often struggle to adequately capture short-range local context [60, 61, 32, 62]. Among the proposed variants, multi-view self-attention (MVA) [32] is particularly noteworthy: each attention head is assigned a distinct temporal window, enabling it to focus on different receptive-field ranges. This design simultaneously enhances the modeling of global dependencies and strengthens the ability to capture local contextual patterns. Such a property aligns well with our goal, as it allows the uncertainty estimation module to effectively capture both long-range dependencies and fine-grained local patterns, which are essential for robust and reliable uncertainty estimation, as shown in Fig. 4. For each attention head, the receptive-field size is defined as:
@@
where $h\in\{0,\ldots,H-1\}$ denotes the head index among a total of $H$ heads, and $w_{h}$ represents the full temporal window size for the $h$-th head.
@@ VIII-A Visualization of Uncertainty Estimation
To verify that the estimated covariance indeed aligns with our objective, we visualize the temporal variation of the precision matrix (i.e., the inverse of the covariance matrix). Specifically, we average the precision matrix along the channel dimension and plot its values over time, as shown in Fig. 5. The input utterance is 60 seconds long, and we apply two types of distortions: adding noise and masking regions of the spectrogram. From the two visualizations, we observe that the precision values decrease as the speech becomes more severely corrupted. Moreover, even regions containing weak or sparse speech exhibit lower precision, indicating reduced reliability. These observations confirm that the estimated frame-level uncertainty effectively reflects the quality and trustworthiness of the speech signal.
@@ VIII-B The Effectiveness of Speaker-level Uncertainty Supervision
We have shown that the frame-level uncertainty estimation within an utterance is accurate. However, because the model applies a softmaxsoftmax function to normalize uncertainty scores within each utterance, the utterance-level uncertainty $\mathbf{\Sigma}^{s}$ is not explicitly supervised or constrained, and may therefore be unreliable. Therefor, we propose speaker-level uncertainty supervision mechanism and verify the effectiveness of it in this section.
As shown in Table I, we compare the performance of our two Stochastic Variance Loss variants, $\mathcal{L}_{\text{SVL-Pre}}$ and $\mathcal{L}_{\text{SVL-Epoch}}$. The results indicate that using $\mathcal{L}_{\text{SVL}}$ alone does not yield noticeable improvements. However, combining it with the uncertainty-aware cosine scoring function leads to better performance. We further evaluate two strategies for the scaling factor $\rho$: (1) the static factor $1/d$ proposed by Wang [38], and (2) a learnable factor $\alpha$ trained jointly with the model. Results demonstrate that the learnable scaling factor significantly outperforms the static counterpart. When combining $\mathcal{L}_{\text{SVL}}$ with the learnable scaling factor $\alpha$, consistent and significant improvements are observed across different loss weights $\lambda$. We further evaluate the models on a cross-domain test set, as shown in Fig. 6. Exps. 3–5 consistently outperform the two baseline methods across all metrics, whereas Exps. 6–8 exhibit unstable behavior and, in some cases, even underperform compared to the baselines. These results indicate that $\mathcal{L}_{\text{SVL-Pre}}$ is more effective, providing stronger robustness and better generalization to unseen domains. Additionally, Exp. 4 with $\lambda=0.05$ achieves the best overall performance.
@@
9 $\mathcal{L}_{\text{AAM}_{u}}$ $\mathbf{I}$ 1.090 1.159 2.188 -4.33% 2.079 2.314 0.91%
@@
10 $\mathcal{L}_{\text{AAM}_{u}}$ $\mathbf{I}-\Delta_{\cos}$ 1.010 1.120 2.107 0.74% 2.199 1.998 4.73%
@@
11 $\mathcal{L}_{\text{AAM}_{u}}$ $\mathbf{I}-\Delta_{\cos}$ 1.128 1.137 2.122 -3.94% 2.128 2.433 -2.79%
@@
12 $\mathcal{L}_{\text{AAM}_{u}}$ $0.5\,\mathbf{I}-\Delta_{\cos}$ 0.995 1.087 1.957 4.53% 2.050 2.214 3.73%
@@ VIII-C The Effectiveness of Global-level Uncertainty Supervision
Speaker-level supervision constrains uncertainty only within utterances from the same speaker and cannot compare uncertainty across different speakers. In practice, however, each training sample produces a different loss value, reflecting its varying difficulty. These loss differences are back-propagated to the uncertainty-aware scale $s_{u}$ in $\mathcal{L}_{\text{AAM}_{u}}$, causing the scale to update dynamically. Since the scale embeds the uncertainty term, this dynamic adjustment naturally induces global-level uncertainty supervision across all samples, enabling the model to learn consistent uncertainty behavior beyond individual speakers. In this section, we explore the effectiveness of the proposed global-level uncertainty supervision.
VIII-C1 Effect of the Scale
As far as we know, few studies have systematically examined the effect of the scale parameter in speaker recognition. In this section, we investigate the effect of the scale $s$ and how varying its value influences the performance of the speaker recognition system. The default scale in AAM-Softmax is $s=32$. Based on this setting, we examine how the performance changes when the scale is increased or decreased. Specifically, we design two additional experiments: (1) a linearly increasing schedule where the scale grows from 32 to 48, and (2) a linearly decreasing schedule where the scale drops from 32 to 16. As shown in Fig. 7, increasing the scale generally improves performance on the hard test set (Vox1-H), while yielding slightly worse results on the easier sets (Vox1-O and Vox1-E). This behavior arises because a larger scale sharpens the softmax distribution, causing the gradient to emphasize harder samples. In contrast, decreasing the scale consistently degrades performance across all test sets, indicating that a small scale severely harms discriminability.
VIII-C2 Ablation Study on the Bias Term $\mathbf{\Lambda}$
@@
Table II reports the performance of our proposed uncertainty-aware AAM-Softmax ($\mathcal{L}_{\text{AAM}_{u}}$ Equation 19). Exp. 9 adopts a static constant Identity matrix as the bias term $\mathbf{\Lambda}$, since the covariance matrix $\mathbf{\Sigma}^{s}\geq 0$, the resulting uncertainty-aware scale $s_{u}$ is always less than 1. Consequently, the final scale applied during training $s\cdot s_{u}$ is consistently reduced. According to our analysis in Section VIII-C1, it is detrimental to performance. As expected, Exp. 9 fails to outperform the baseline.
Exp. 10 adopts dynamically varying $\mathbf{\Lambda}$ derived from $\Delta_{\text{cos}}=(\cos\theta_{y_{i}}-\displaystyle\max_{j\neq y_{i}}\cos\theta_{j}).\texttt{detach()}$, i.e., the difference between the target cosine and the maximum non-target cosine, so that $\Delta_{\text{cos}}$ is not influenced by parameter updates during training and is only used as a difficulty indicator. This term can be interpreted as a measure of how difficult a sample is to learn: a larger $\Delta_{\text{cos}}$ implies that the sample is easier. Empirically, we find that most $\Delta_{\text{cos}}$ values lie within the range of –0.8 to 0.5.
TABLE IV: Ablation study of progressively incorporating each component in the proposed $\mathcal{U}^{3}$-xi model. ECAPA-TDNN is used as the speaker encoder in all experiments.

### v2

@@ V-A Uncertainty-aware Softmax
Conventional Softmax-based objectives, as in (7) (8), operate solely on the embedding $\boldsymbol{\phi}^{\text{s}}$, which represents only a point estimate of the speaker identity. Such formulations overlook the covariance information $\mathbf{\Sigma}^{\text{s}}$, thereby discarding valuable information about embedding reliability.
@@
To address this limitation, we reformulate the $l_{2}$ normalization by incorporating the uncertainty. The uncertainty-aware normalized embedding is given by:
@@
$\tilde{\boldsymbol{\phi}}^{s}$ $=\frac{\boldsymbol{\phi}^{\text{s}}}{\|\boldsymbol{\phi}^{\text{s}}\|}\cdot\frac{\sqrt{{\boldsymbol{\phi}^{\text{s}}}^{\top}(\mathbf{\Lambda}+\mathbf{\Sigma}^{\text{s}})^{-1}\boldsymbol{\phi}^{\text{s}}}}{\|\boldsymbol{\phi}^{\text{s}}\|},$ (14)
which normalizes the embedding under the Mahalanobis metric induced by the uncertainty covariance $\mathbf{\Lambda}+\mathbf{\Sigma}^{\text{s}}$ where $\mathbf{\Lambda}$ is a bias term that enhances numerical stability to avoid small denominators. This formulation ensures that $\tilde{\boldsymbol{\phi}}^{s}$ is a unit vector in the uncertainty-aware space, reflecting both its magnitude and its uncertainty structure. However, in practice, explicitly computing the inverse $(\mathbf{\Lambda}+\mathbf{\Sigma}^{\text{s}})^{-1}$ often leads to unstable gradients during training. To mitigate this issue, we adopt a numerically stable alternative:
$\tilde{\boldsymbol{\phi}}^{s}$ $=\frac{\boldsymbol{\phi}^{\text{s}}}{\|\boldsymbol{\phi}^{\text{s}}\|}\cdot\frac{\|\boldsymbol{\phi}^{\text{s}}\|}{\sqrt{{\boldsymbol{\phi}^{\text{s}}}^{\top}(\mathbf{\Lambda}+\mathbf{\Sigma}^{\text{s}})\boldsymbol{\phi}^{\text{s}}}}.$ (15)
These two expressions are closely related. We visualize the value of the second multiplicative term (i.e., the factor after $\cdot$”) in Fig. 2, where both the bias term $\mathbf{\Lambda}$ and the embeddings $\boldsymbol{\phi}^{s}$ are drawn from a Gaussian distribution. The figure shows that the two formulations exhibit similar trends, while differing mainly in their absolute values. Moreover, as the mean variance increases, the discrepancy between the two terms also becomes larger.
Figure 2: The value difference of the second multiplicative term.
Furthermore, we rewrite Equation 15 as
$\tilde{\boldsymbol{\phi}}^{s}$ $=\frac{\boldsymbol{\phi}^{\text{s}}}{\|\boldsymbol{\phi}^{\text{s}}\|}\cdot s_{u},$ (16)
where $s_{u}>0$ is an uncertainty-aware scale that summarizes the effect of the second multiplicative term. Importantly, $s_{u}$ increases as the uncertainty decreases, thereby modulating the embedding magnitude according to the confidence of its estimate.
@@
We then incorporate the uncertainty-aware embedding $\tilde{\boldsymbol{\phi}}^{s}$ into the AAM-Softmax formulation to obtain the uncertainty AAM-Softmax (UAAM):
$\mathcal{L}_{\text{UAAM}}=$ $-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{s\cdot s_{u}\cos(\theta_{y_{i}}+m)}}{e^{s\cdot s_{u}\cos(\theta_{y_{i}}+m)}+\sum_{j=1,j\neq y_{i}}^{C}e^{s\cdot s_{u}\cos\theta_{j}}}.$ (17)
@@ V-B Gradient-Driven Adaptation of Scale and Covariance
From a training perspective, the uncertainty-aware scale $s_{u}$ enables the model to adaptively modulate the loss according to the reliability of each embedding. Since gradient-based optimization always aims to minimize the training loss, the direction in which $s_{u}$ is adjusted depends on how loss reduction can be most effectively achieved for different samples.
For unreliable or incorrectly classified samples, which typically incur large loss values, reducing the scale $s_{u}$ provides a natural descent direction. A smaller scale smooths the softmax distribution, attenuating the influence of incorrect logits and thereby reducing the loss. In our formulation, this decrease in $s_{u}$ directly corresponds to an increase in the predicted variance (uncertainty), reflecting the higher ambiguity of hard samples.
Conversely, for easy and reliable samples that are already correctly classified and incur small losses, the optimizer still seeks further loss minimization. In this regime, increasing $s_{u}$ sharpens the softmax distribution and amplifies the target-class logit, pushing the predicted posterior probability closer to one. Under the cross-entropy loss, this leads to additional loss reduction, and thus the back-propagation gradient naturally drives $s_{u}$ upward. As a result, the predicted variance $\mathbf{\Sigma}^{\text{s}}$ decreases, indicating higher confidence in these embeddings.
In summary, hard samples are associated with smaller scales and larger variances, while easy samples yield larger scales and lower variances. This scale–variance coupling allows uncertainty to be learned through gradient-driven loss minimization, as shown in Fig. 3.
Hard Sample (Large Loss)Gradient BackpropagationDecrease $s_{u}$Soften Softmax DistributionIncrease $\mathbf{\Sigma}^{\text{s}}$Easy Sample (Small Loss)Gradient BackpropagationIncrease $s_{u}$Sharpen Softmax DistributionDecrease $\mathbf{\Sigma}^{\text{s}}$
Figure 3: Illustration of gradient-driven scale adaptation.
@@ VI Uncertainty Estimation Module
The previous two sections focus on modifying the training strategies without altering the model architecture. In this section, we aim to redesign the uncertainty estimation module. Our initial work xi-vector [24], the uncertainty estimation relies on two simple learnable layers with basic activation functions to predict frame-level uncertainty, as illustrated in Fig. 1. However, this design is relatively shallow and lacks the ability to capture temporal dependencies across speech frames. Our previous work xi+ [36] addressed this limitation by incorporating a Transformer encoder to model temporal relations, thereby enhancing the model’s capacity to capture complex non-linear patterns. Although Transformer-based architectures have been explored for speaker recognition, they often struggle to adequately capture short-range local context [60, 61, 32, 62]. Among the proposed variants, multi-view self-attention (MVA) [32] is particularly noteworthy: each attention head is assigned a distinct temporal window, enabling it to focus on different receptive-field ranges. Such a property aligns well with our goal, as it allows the uncertainty estimation module to effectively capture both long-range dependencies and fine-grained local patterns, which are essential for robust and reliable uncertainty estimation, as shown in Fig. 4. For each attention head, the receptive-field size is defined as:
@@
where $h\in\{0,\ldots,H-1\}$ denotes the head index among a total of $H$ heads, and $w_{h}$ represents the temporal window size for the $h$-th head.
@@ VIII-A Visualization of Uncertainty Estimation
To verify that the estimated covariance indeed aligns with our objective, we visualize the temporal variation of the precision matrix (i.e., the inverse of the covariance matrix). Specifically, we average the precision matrix along the channel dimension and plot its values over time, as shown in Fig. 5. The input utterance is 60 seconds long, and we apply two types of distortions: adding noise and masking partial regions of the spectrogram. From the two visualizations, we observe that the precision values decrease as the speech becomes severely corrupted. Moreover, even regions containing weak or sparse speech exhibit lower precision, indicating reduced reliability. These observations confirm that the estimated frame-level uncertainty effectively reflects the quality and trustworthiness of the speech signal.
@@ VIII-B The Effectiveness of Speaker-level Uncertainty Supervision
We have shown that the frame-level uncertainty estimation within an utterance is accurate. However, the model applies a softmaxsoftmax function to normalize uncertainty scores within each utterance, which preserves only the internal ordering of frame-wise uncertainty while eliminating the absolute scale. As a result, uncertainty values across different utterances become incomparable, and $\mathbf{\Sigma}^{\text{s}}$ cannot be reliably interpreted or supervised at the utterance level. To address this issue, we propose a speaker-level uncertainty supervision mechanism and verify its effectiveness in this section.
As shown in Table I, we compare the performance of our two Stochastic Variance Loss variants, $\mathcal{L}_{\text{SVL-Pre}}$ and $\mathcal{L}_{\text{SVL-Epoch}}$. The results indicate that using $\mathcal{L}_{\text{SVL}}$ alone does not yield noticeable improvements. However, combining it with the uncertainty-aware cosine scoring strategy leads to better performance. We further evaluate two strategies for the scaling factor $\rho$: (1) the static factor $1/d$ proposed by Wang [38], and (2) a learnable factor $\alpha$ trained jointly with the model proposed by our previous work xi+ [36]. Results demonstrate that the learnable scaling factor significantly outperforms the static counterpart. When combining $\mathcal{L}_{\text{SVL}}$ with the learnable scaling factor $\alpha$, consistent and significant improvements are observed across different loss weights $\lambda$. We further evaluate the models on a cross-domain test set, as shown in Fig. 6. Exps. 3–5 consistently outperform the two baseline methods across all metrics, whereas Exps. 6–8 exhibit unstable behavior and, in some cases, even underperform compared to the baselines. These results indicate that $\mathcal{L}_{\text{SVL-Pre}}$ is more effective, providing stronger robustness and better generalization to cross-domains. Additionally, Exp. 4 with $\lambda=0.05$ achieves the best overall performance.
@@
9 $\mathcal{L}_{\text{UAAM}}$ $\mathbf{I}$ 1.090 1.159 2.188 -4.33% 2.079 2.314 0.91%
@@
10 $\mathcal{L}_{\text{UAAM}}$ $\mathbf{I}-\Delta_{\cos}$ 1.010 1.120 2.107 0.74% 2.199 1.998 4.73%
@@
11 $\mathcal{L}_{\text{UAAM}}$ $\mathbf{I}-\Delta_{\cos}$ 1.128 1.137 2.122 -3.94% 2.128 2.433 -2.79%
@@
12 $\mathcal{L}_{\text{UAAM}}$ $0.5\,\mathbf{I}-\Delta_{\cos}$ 0.995 1.087 1.957 4.53% 2.050 2.214 3.73%
@@ VIII-C The Effectiveness of Global-level Uncertainty Supervision
Speaker-level supervision constrains uncertainty only within utterances from the same speaker and cannot compare uncertainty across different speakers. In practice, however, each training sample produces a different loss value, reflecting its varying difficulty. These loss differences are back-propagated to the uncertainty-aware scale $s_{u}$ in $\mathcal{L}_{\text{UAAM}}$, causing the scale to update dynamically. In this section, we explore the effectiveness of the proposed global-level uncertainty supervision.
@@
VIII-C1 Preliminary Analysis of Scale Effects
As far as we know, few studies have systematically examined the effect of the scale parameter in speaker recognition. In this section, we investigate the effect of the scale $s$ and how varying its value influences the performance of the speaker recognition system. The default scale in AAM-Softmax is $s=32$. Based on this setting, we examine how the performance changes when the scale is increased or decreased. Specifically, we design two additional experiments: (1) a linearly increasing schedule where the scale grows from 32 to 48, and (2) a linearly decreasing schedule where the scale drops from 32 to 16. As shown in Fig. 7, increasing the scale generally improves performance on the hard test set (Vox1-H), while yielding slightly worse results on the easier sets (Vox1-O and Vox1-E). This behavior arises because a larger scale sharpens the softmax distribution, causing the gradient to emphasize harder samples. In contrast, decreasing the scale consistently degrades performance across all test sets, indicating that decreasing scale severely harms discriminability.
VIII-C2 Ablation Study on the Bias Term $\mathbf{\Lambda}$
Table II reports the performance of our proposed uncertainty-aware AAM-Softmax ($\mathcal{L}_{\text{UAAM}}$ Equation 17). Exp. 9 adopts a static constant Identity matrix as the bias term $\mathbf{\Lambda}$, since the covariance matrix $\mathbf{\Sigma}^{\text{s}}\geq 0$, the resulting uncertainty-aware scale $s_{u}$ is always less than 1. Consequently, the final scale applied during training $s\cdot s_{u}$ is less than the original scale $s$. According to our analysis in the previous section, it is detrimental to performance. As expected, Exp. 9 fails to outperform the baseline.
Exp. 10 adopts a dynamically varying bias term $\mathbf{\Lambda}$ derived from $\Delta_{\text{cos}}=\big(\cos\theta_{y_{i}}-\max_{j\neq y_{i}}\cos\theta_{j}\big).\texttt{detach()}$, which measures the cosine margin between the target class and the most competitive non-target class. The detach() operation prevents gradients from flowing through $\Delta_{\text{cos}}$, ensuring that it serves solely as a difficulty indicator rather than a trainable quantity. Intuitively, $\Delta_{\text{cos}}$ reflects the sample-level classification difficulty: larger values indicate that the target class is already well separated from non-target classes, corresponding to easier samples, whereas smaller or negative values imply harder samples. Empirically, we observe that most $\Delta_{\text{cos}}$ values fall within the range of [0.8, 0.5][-0.8,\,0.5].
TABLE IV: Ablation study of progressively incorporating each component in the proposed $\mathcal{U}^{3}$-xi model.